OpenStudy (anonymous):

in the xy plane the graph of the function is a line.If h(-1)=4 and h(5)=1 .What is the value of h(0) ?

OpenStudy (anonymous):

@calculusfunctions help

OpenStudy (anonymous):

@zordoloom

OpenStudy (anonymous):

Can you form an equation from this?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

@zordoloom

OpenStudy (anonymous):

okay, so one point would be (-1, 4). Can you find the other point?

OpenStudy (anonymous):

(5,1)

OpenStudy (campbell_st):

the value of h(0) will be the y- intercept of the line.

OpenStudy (anonymous):

Good, now use these values to find the slope.

OpenStudy (anonymous):

how @campbell_st ?

OpenStudy (calculusfunctions):

@girlish do you still want my help on this?

OpenStudy (anonymous):

slope is -1/2

OpenStudy (anonymous):

Thats pretty close. But not exactly right.

OpenStudy (campbell_st):

well you have ordered pairs (1, 4) and (5, 1) find the slope the equation of the line will be h(x) =mx + b m = slope of the line

OpenStudy (anonymous):

yes what's the next step? @calculusfunctions

OpenStudy (campbell_st):

actually I think that is the slope

OpenStudy (anonymous):

After You find the slope (-2/3). Use the point slope formula and one of the points.

OpenStudy (campbell_st):

oops (-1, 4) and (5, 1)

OpenStudy (campbell_st):

how do you get - 2/3 ...?

OpenStudy (campbell_st):

(1 - 4)/(5 - -1) = -3/6 = -1/2

OpenStudy (anonymous):

oh, I thought it was (5,0) my bad.

OpenStudy (anonymous):

yes it's -1/2

OpenStudy (anonymous):

-1/2 is correct.

OpenStudy (campbell_st):

@girlish just substitute one of the ordered pairs into \[y = -\frac{1}{2} x + b\] and solve for b the value of h(0) will be b

OpenStudy (anonymous):

options for the question are 2.2 3.3 3.5 3.7

OpenStudy (anonymous):

Wouldn't that be too easy? Just plug in the slope and one of the points into the point slope formula and solve for y. You'll have your y intercept, which is what you are looking for.

OpenStudy (anonymous):

3.5 would be the aanswer.right?

OpenStudy (anonymous):

yes, that would be correct.

OpenStudy (campbell_st):

thats all you need to do... you can use the slope point formula... but will get the same answer

OpenStudy (campbell_st):

that it... well done

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yep.

OpenStudy (calculusfunctions):

@girlish h(-1) = 4 is the same as (-1, 4) and h(5) = 1 is the same as (5, 1). Thus the slope as you said correctly is -1/2. Now use this slope as the m and one of the given points as (x, y) to solve for b in the equation y =mx + b. I see you already got it.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

can u simplify (x+2)(x-2)(x^4+16)(x^2+4)

OpenStudy (anonymous):

@calculusfunctions

OpenStudy (anonymous):

i solved part of it (x+2)(x-2) will become x^2-4

OpenStudy (calculusfunctions):

Now multiply (x² - 4)(x² + 4) in a similar manner.

OpenStudy (anonymous):

{x^4 +(2)^4}+{x^2+(2)^2}

OpenStudy (anonymous):

you told me to expand

OpenStudy (anonymous):

how did it becum (x² - 4)(x² + 4)?

OpenStudy (anonymous):

be quick

OpenStudy (calculusfunctions):

When you multiply two binomials, that is expanding. I am asking you to multiply (x² - 4)(x² + 4)

OpenStudy (anonymous):

x^4-16

OpenStudy (calculusfunctions):

Don't rush me if you want my help. Yes!! Now multiply this result by the remaining binomial (x^4 + 16)

OpenStudy (anonymous):

the answer would be x^8-256 .Right?

OpenStudy (calculusfunctions):

YES!

OpenStudy (calculusfunctions):

That's the final answer!

OpenStudy (anonymous):

:)